OpenStudy (anonymous):

help plzzzzz the x-intercept of x + 2 = -3(y - 2)2?

OpenStudy (lgbasallote):

x-intercept -> value of x when y = 0 so change all y to zero then solve for x..get the idea? :D

OpenStudy (anonymous):

yep. just plug in a zero for y and solve for x....

OpenStudy (anonymous):

34?

OpenStudy (lgbasallote):

hmmm... you let y = 0 right? x+2 = -3 (0-2) 2

OpenStudy (anonymous):

yess

OpenStudy (lgbasallote):

i guess you can solve from there? hint: it's not 34 :DD

OpenStudy (anonymous):

x+2 = -3 (0-2)^ 2

OpenStudy (lgbasallote):

ohhh...okay then...try to solve from there :DD hint: it's still not 34 :p hehe

OpenStudy (anonymous):

-9?

OpenStudy (anonymous):

-9?

OpenStudy (lgbasallote):

you might want to recheck your solution..some sign problems maybe?

OpenStudy (anonymous):

oh oh yeah didnt square the 2 properly, i get -1 now

OpenStudy (lgbasallote):

-1? still no...be careful in solving ^_^ remember...when you square a negative, the answer is positive. positive times negative is negative.

OpenStudy (anonymous):

ok so, x+2=-3+4 x= .3+4-2

OpenStudy (lgbasallote):

+? on your problem it states 3(-2)^2..that's multiplication

OpenStudy (anonymous):

oh my god your right!i wrote the problem down wrong on my paper!

OpenStudy (lgbasallote):

hahaha =)))) so now you get your mistake :p

OpenStudy (anonymous):

yes, thankyou hahah

OpenStudy (lgbasallote):

<tips hat>